# Adversarial bandit optimization for approximately linear functions

### v5

@@ 1 Introduction
For convex loss functions, the bandit optimization has been extensively studied (see, e.g.,[3, 9, 17]). $\mathcal{O}(d^{1/3}T^{3/4})$ regret bounds are shown by Flaxman et al. [10]. Lattimore [16] shows an information-theoretic regret bound $\widetilde{O}(d^{2.5}\sqrt{T})$ for convex loss functions. For linear loss functions, Abernethy et al. [3] propose the SCRiBLe algorithm and give an expected regret bound $\mathcal{O}(d\sqrt{T\ln T})$, achieving optimal dependence on $T$ [8]. Lee et al. [17] propose SCRiBLe with lifting and increasing learning rates and show a high-probability regret bound $\widetilde{O}(d^{2}\sqrt{T})$.
Recently, non-convex functions are also getting popular in this literature. For example, Agarwal et al. [4] show a regret bound $\mathcal{O}(poly(d)T^{2/3})$ for smooth and bounded non-convex functions. Ghai et al. [12] propose algorithms with regret bounds $\mathcal{O}(poly(d)T^{2/3})$ under the assumption that non-convex functions are reparameterized as some convex functions.
In this paper, we investigate the bandit optimization problem for a class of non-convex and non-smooth loss functions. The function class consists of non-smooth and non-convex functions that are "close" to linear functions, in the sense that functions in the class can be viewed as linear functions with adversarial non-convex perturbations whose amount is up to $\epsilon$. Bandit optimization for linear loss functions with stochastic noise (e.g., [1, 5]) cannot be applied to our problem. Moreover, standard bandit linear optimization methods for gradient estimation, such as those based on self-concordant barrier regularizers [13], cannot be effectively applied to our problem. We propose a novel approach to analyzing high-probability regret, introducing a new method for decomposing regret.
@@
    When $\epsilon\neq 0$, we propose a modification of SCRiBLe with lifting and increasing learning rates [17] and leverage the properties of the $\nu$-normal barrier [19] to prove its high probability regret bound $\widetilde{O}(d\sqrt{T}+\epsilon dT)$. We also derive its expected regret $\mathcal{O}(d\sqrt{T\ln T}+\epsilon dT)$.
@@
    When $\epsilon=0$, this problem becomes bandit linear optimization, a special case with no perturbation. Compared to the result of [17], holding with probability $1-\gamma$, $\mathcal{O}(\ln^{2}(dT)d^{2}\ln T\sqrt{T\ln\frac{\ln(dT)}{\gamma}})$, we use a different regret decomposition approach to achieve a better high-probability regret bound $\mathcal{O}(d\sqrt{T\ln T}+\ln T\sqrt{T\ln(\frac{\ln T}{\gamma})}+\ln(\frac{\ln T}{\gamma}))$.
@@
    We prove a lower bound $\Omega(\epsilon T)$, implying that our bounds are tight w.r.t. the parameter $\epsilon$.
@@ 2 Related Work
Bandit linear optimization was first proposed by Awerbuch & Kleinberg [6], who achieved a regret bound of $\mathcal{O}(d^{3/5}T^{2/3})$ against oblivious adversaries. Later, McMahan & Blum [18] established a regret bound of $\mathcal{O}(dT^{3/4})$ when facing adaptive adversaries. A foundational approach in bandit optimization problems involves gradient-based smoothing techniques. Flaxman et al. [10] proposed a method for constructing an unbiased estimator of the loss function’s gradient in the bandit setting. Abernethy et al. [3] introduced the SCRiBLe algorithm and achieved an expected regret bound of $\mathcal{O}(d\sqrt{T\ln T})$ against oblivious adversaries. Bartlett et al. [7] proposed a high-probability regret bound of $\mathcal{O}(d^{2/3}\sqrt{T\ln dT})$ under a special condition. Subsequently, Lee et al. [17] presented a high-probability regret bound $\widetilde{O}(d^{2}\sqrt{T})$ for both oblivious and adaptive adversaries. In recent work, Ito and Takemura [14, 15] proposed a bandit linear algorithm that adapts to both stochastic and adversarial environments, achieving a regret bound of $\mathcal{O}(d\sqrt{T\ln T})$ in the adversarial setting. Rodemann et al. [21] established a connection between bandit problems and Bayesian black-box optimization, offering theoretical foundations for regret bounds across both domains.
Unlike bandit convex optimization problems, which have been extensively explored and analyzed, bandit non-convex optimization problems introduce unique challenges due to the complexity of exploring and exploiting in a non-convex area. Gao et al. [11] considered both non-convex losses and non-stationary data and established a regret bound of $O(\sqrt{T+poly(T)}$. Yang et al. [22] achieved a regret bound of $O(\sqrt{T\log T})$ for non-convex loss functions. However, they both require the loss functions to have smoothness properties, and our loss functions are neither convex nor smooth.
2.1 Comparison to Lee et al. [17]
@@
    Revised Generality of Problem Setting: Like the SCRiBLe algorithm [3], our approach is more general, treating bandit linear optimization as a special case within a broader problem framework.
@@ 3.1 Notation
[…]
For any $v\in\mathbb{R}^{d}$, let $v^{\perp}$ be the space orthogonal to $v$. Let $\mathbb{S}^{d}_{1}=\{x\mid\|x\|=1\}$. The vector $e_{i}\in\mathbb{R}^{d}$ is a standard basis vector with a value of 111 in the $i$-th position and 0 in all other positions. $I$ is an identity matrix with dimensionality implied by context.
@@ 3.2 Problem Setting
Let $\mathcal{K}\subseteq\mathbb{R}^{d}$ be a bounded and closed convex set such that for any $x,y\in\mathcal{K}$, $\|x-y\|\leq D$. Furthermore, we assume that $\mathcal{K}$ contains the unit ball centered at the zero vector. Otherwise, we can apply an affine transformation to translate the center point of the convex set to the origin. Let $\mathcal{K^{\prime}}=\{(x,1):x\in\mathcal{K}\}$. For any $\delta\in(0,1)$, let $\mathcal{K}_{\delta}=\{x|\frac{1}{1-\delta}x\in\mathcal{K}\}$ and $\mathcal{K^{\prime}_{\delta}}=\{(x,1):x\in\mathcal{K_{\delta}}\}$, respectively.
@@
In this paper, we consider the bandit optimization $(\mathcal{K},\mathcal{F})$, where $\mathcal{F}$ is the set of $\epsilon$-approximately linear functions $f(x)=\theta^{\top}x+\sigma(x)$ with $\|\theta\|\leq G$. Bandit optimization for $\epsilon$-approximately linear functions can be defined as the following statement. For every round $t=1,..,T$, the player selects an action $x_{t}\in\mathcal{K}$, without knowing the loss function in advance. The environment, modeled as an oblivious adversary, chooses a sequence of linear loss vectors $\theta_{1},\theta_{2},...,\theta_{t}$ before the interaction begins and independent of the player’s actions. After selecting $x_{t}$, the adversary chooses a perturbation $\sigma_{t}(x_{t})(|\sigma_{t}(x_{t})|\leq\epsilon)$ and the player observes only the scalar loss $f_{t}(x_{t})(=\theta^{\top}_{t}x_{t}+\sigma_{t}(x_{t}))$. The goal of the player is to minimize the regret $\sum_{t=1}^{T}f_{t}(x_{t})-\min_{x\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(x)$.
@@ 4 Main Results
[…]
2:Initialize: $x^{\prime}_{1}=\arg\min_{x^{\prime}\in\mathcal{K^{\prime}}}{\mathcal{R}(x^{\prime})}$
@@
7:  Update $x^{\prime}_{t+1}=\arg\min\limits_{x^{\prime}\in\mathcal{K^{\prime}}}{\eta\sum_{\tau=1}^{t}g_{\tau}^{\top}x^{\prime}+\mathcal{R}(x^{\prime})}$
@@ 4.1 SCRiBLe with lifting
For the decision set $\mathcal{K}$ with a $\nu$-normal barrier on $con(\mathcal{K})$, where $con(\mathcal{K})=\{\mathbf{0}\}\cup\{(x,b):\frac{x}{b}\in\mathcal{K},x\in\mathbb{R}^{d},b>0\}$, we apply Algorithm 1 to $\epsilon$-approximately linear functions. Recall $\mathcal{K^{\prime}}=\{(x,1):x\in\mathcal{K}\}$.
We simplify SCRiBLe with lifting and increasing learning rates [17]. The algorithm performs original SCRiBLe [3] over a lifted decision set, using a $\nu$-normal barrier $\mathcal{R}$ defined over the $con(\mathcal{K})$ (which always exists) as the regularizer to generate the sequence $x^{\prime}_{1},...,x^{\prime}_{t}$. It set $y^{\prime}_{t}=x^{\prime}_{t}+\mathbf{A}_{t}\mu_{t}$, where $\mathbf{A}_{t}=[\nabla^{2}\mathcal{R}(x^{\prime}_{t})]^{-\frac{1}{2}}$ and $\mu_{t}$ is uniformly sampled at random from $\mathbb{S}_{1}^{d+1}\cap(\mathbf{A}_{t}e_{d+1})^{\perp}$. Since $\mu_{t}$ is orthogonal to $\mathbf{A}_{t}e_{d+1}$, the last coordinate of $\mathbf{A}_{t}\mu_{t}$ is zero, ensuring that $y^{\prime}_{t}=(y_{t},1)$ remains within $\mathcal{K^{\prime}}$. The actual point played is still $y_{t}$. After playing $y_{t}$ and observing $f_{t}(y_{t})(=\theta_{t}^{\top}y_{t}+\sigma_{t}(y_{t}))$, it constructs the loss estimator the same way as SCRiBLe algorithm [3]: $g_{t}=df_{t}(y_{t})\mathbf{A}_{t}^{-1}\mu_{t}$. Furthermore, it adopts the same update method as FTRL algorithm [13]: $x^{\prime}_{t+1}=\arg\min\limits_{x^{\prime}\in\mathcal{K^{\prime}}}{\eta\sum_{\tau=1}^{t}g_{\tau}^{\top}x^{\prime}+\mathcal{R}(x^{\prime})}$.
Compared with SCRiBLe with lifting and increasing learning rates [17], which focuses on bandit linear optimization, we do not use the increasing learning rates part, but retain the lifting. This preserves its advantages; for instance, the $\nu$-normal barrier $\mathcal{R}$ always exists on $con(\mathcal{K})$ and we can leverage the properties of the $\nu$-normal barrier. When the loss takes the form $f_{t}(y_{t})=\theta_{t}^{\top}y_{t}$ (i.e., when $\epsilon=0$), the analysis by Lee et al. [17] shows that the first $d$ coordinates of $g_{t}$ are indeed an unbiased estimator of $\theta_{t}$. However, when $\epsilon\neq 0$ since $g_{t}$ will always be influenced by $\sigma_{t}(y_{t})$, $g_{t}$ is no longer an unbiased estimator of $\theta_{t}$.
Assumption 1
Let $x^{*}=\arg\min_{x\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(x)$, $x_{\delta}^{*}=\Pi_{\mathcal{K}_{\delta}}(x^{*})$, and $x_{\delta}^{*^{\prime}}=(x_{\delta}^{*},1)$. Assume that for a $\nu$-normal barrier $\mathcal{R}$ on $con(\mathcal{K})$, the following inequality holds: $\nabla\mathcal{R}(x^{\prime}_{t})^{\top}(x_{\delta}^{*^{\prime}}-x^{\prime}_{t})\geq-\nu$.
Assumption 1 is introduced to enable the use of Lemma 4 in proving Lemma 8. Under this assumption, the theoretical upper bound can be established. However, when $\nabla\mathcal{R}(x^{\prime}_{t})^{\top}(x_{\delta}^{*^{\prime}}-x^{\prime}_{t})<-\nu$, our method no longer admits the theoretical upper bound.
@@
Assume Assumptions 1. The algorithm with parameters $\eta=\frac{\sqrt{2\nu\log T}}{2d\sqrt{T}}$, $\delta=\frac{1}{T^{2}}$ guarantees the following expected regret bound
$\mathbb{E}[\sum_{t=1}^{T}f_{t}(y_{t})-\min_{x\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(x)]\leq 4d\sqrt{2\nu T\log T}+\frac{GD}{T}+2T\epsilon+dT\epsilon(2\nu+\sqrt{\nu}).$ (2)
@@
Assume Assumptions 1. The algorithm with parameters $\eta=\frac{\sqrt{2\nu\ln T}}{2d\sqrt{T}}$, $\delta=\frac{1}{T^{2}}$ ensures that with probability at least $1-\gamma$
$\begin{array}[]{l}\sum_{t=1}^{T}f_{t}(y_{t})-\min_{x\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(x)\leq 4d\sqrt{2\nu T\ln T}+\frac{GD}{T}+2GD\ln\frac{C}{\gamma}+2T\epsilon\\ +C(1+\epsilon)\sqrt{8T\ln\frac{C}{\gamma}}+Td\epsilon(2\nu+\sqrt{\nu})\end{array}$

### v8

@@ 1 Introduction
For convex loss functions, the bandit optimization has been extensively studied (see, e.g.,[abernethy2012interior, dani2007price, lee2020bias]). $\mathcal{O}(d^{1/3}T^{3/4})$ regret bounds are shown by Flaxman et al. [flaxman-etal:soda05]. Lattimore [lattimore:msl20] shows an information-theoretic regret bound $\widetilde{O}(d^{2.5}\sqrt{T})$ for convex loss functions. For linear loss functions, Abernethy et al. [abernethy2012interior] propose the SCRiBLe algorithm and give an expected regret bound $\mathcal{O}(d\sqrt{T\ln T})$, achieving optimal dependence on $T$ [bubeck2012towards]. Lee et al. [lee2020bias] propose SCRiBLe with lifting and increasing learning rates and show a high-probability regret bound $\widetilde{O}(d^{2}\sqrt{T})$.
Recently, non-convex functions are also getting popular in this literature. For example, Agarwal et al. [agarwal2019learning] show a regret bound $\mathcal{O}(poly(d)T^{2/3})$ for smooth and bounded non-convex functions. Ghai et al. [ghai2022non] propose algorithms with regret bounds $\mathcal{O}(poly(d)T^{2/3})$ under the assumption that non-convex functions are reparameterized as some convex functions.
In this paper, we investigate the bandit optimization problem for a class of non-convex and non-smooth loss functions. The function class consists of non-smooth and non-convex functions that are "close" to linear functions, in the sense that functions in the class can be viewed as linear functions with adversarial non-convex perturbations whose amount is up to $\epsilon$. Bandit optimization for linear loss functions with stochastic noise (e.g., [abbasi2011improved, amani2019linear]) cannot be applied to our problem. We propose a novel approach to analyzing high-probability regret, introducing a new method for decomposing regret.
@@
    When $\epsilon\neq 0$, we propose a modification of SCRiBLe with lifting and increasing learning rates [lee2020bias] and leverage the properties of the $\nu$-normal barrier [nemirovski2004interior] to prove its high probability regret bound $\widetilde{O}(d\sqrt{T}+\sqrt{\epsilon}dT+\epsilon dT)$. We also derive its expected regret $\widetilde{O}(d\sqrt{T}+\sqrt{\epsilon}dT+\epsilon dT)$.
@@
    When $\epsilon=0$, this problem becomes bandit linear optimization, a special case with no perturbation. Compared to the result of [lee2020bias], holding with probability $1-\gamma$, $\mathcal{O}(\ln^{2}(dT)d^{2}\ln T\sqrt{T\ln\frac{\ln(dT)}{\gamma}})$, we use a different regret decomposition approach to achieve a better high-probability regret bound $\mathcal{O}(d\sqrt{T\ln T}+\ln T\sqrt{T\ln(\frac{\ln T}{\gamma})}+\ln(\frac{\ln T}{\gamma}))$.
@@
    We prove a lower bound $\Omega(\epsilon T)$, characterizing the minimal dependence on the parameter $\epsilon$.
@@ 2 Related Work
Bandit linear optimization was first proposed by Awerbuch & Kleinberg [awerbuch2004adaptive], who achieved a regret bound of $\mathcal{O}(d^{3/5}T^{2/3})$ against oblivious adversaries. Later, McMahan & Blum [mcmahan2004online] established a regret bound of $\mathcal{O}(dT^{3/4})$ when facing adaptive adversaries. A foundational approach in bandit optimization problems involves gradient-based smoothing techniques. Flaxman et al. [flaxman-etal:soda05] proposed a method for constructing an unbiased estimator of the loss function’s gradient in the bandit setting. Abernethy et al. [abernethy2012interior] introduced the SCRiBLe algorithm and achieved an expected regret bound of $\mathcal{O}(d\sqrt{T\ln T})$ against oblivious adversaries. Bartlett et al. [bartlett2008high] proposed a high-probability regret bound of $\mathcal{O}(d^{2/3}\sqrt{T\ln dT})$ under a special condition. Subsequently, Lee et al. [lee2020bias] presented a high-probability regret bound $\widetilde{O}(d^{2}\sqrt{T})$ for both oblivious and adaptive adversaries. In recent work, Ito and Takemura [ito2023best, ito2023exploration] proposed a bandit linear algorithm that adapts to both stochastic and adversarial environments, achieving a regret bound of $\mathcal{O}(d\sqrt{T\ln T})$ in the adversarial setting. Rodemann et al. [rodemann2024reciprocal] established a connection between bandit problems and Bayesian black-box optimization, offering theoretical foundations for regret bounds across both domains.
Unlike bandit convex optimization problems, which have been extensively explored and analyzed, bandit non-convex optimization problems introduce unique challenges due to the complexity of exploring and exploiting in a non-convex area. Gao et al. [gao2018online] considered both non-convex losses and non-stationary data and established a regret bound of $O(\sqrt{T+poly(T)}$. Yang et al. [yang2018optimal] achieved a regret bound of $O(\sqrt{T\ln T})$ for non-convex loss functions. However, they both require the loss functions to have smoothness properties, and our loss functions are neither convex nor smooth.
2.1 Comparison to Lee et al. [lee2020bias]
@@
    Revised Generality of Problem Setting: Like the SCRiBLe algorithm [abernethy2012interior], our approach is more general, treating bandit linear optimization as a special case within a broader problem framework.
@@ 3.1 Notation
We abbreviate the $2$-norm $\|\cdot\|_{2}$ as $\|\cdot\|$. For a twice differentiable convex function $\mathcal{R}:\mathbb{R}^{d}\to\mathbb{R}$ and any $x,h\in\mathbb{R}^{d}$, let $\|h\|_{x}=\|h\|_{\nabla^{2}\mathcal{R}(x)}=\sqrt{h^{\top}\nabla^{2}\mathcal{R}(x)h}$, and $\|h\|_{x}^{*}=\|h\|_{(\nabla^{2}\mathcal{R}(x))^{-1}}=\sqrt{h^{\top}(\nabla^{2}\mathcal{R}(x))^{-1}h}$, respectively.
For any $v\in\mathbb{R}^{d}$, let $v^{\perp}$ be the space orthogonal to $v$. Let $\mathbb{S}^{d}_{1}=\{x\mid\|x\|=1\}$. The vector $e_{i}\in\mathbb{R}^{d}$ is a standard basis vector with a value of $1$ in the $i$-th position and 0 in all other positions. $I$ is an identity matrix with dimensionality implied by context.
@@ 3.2 Problem Setting
Let $\mathcal{K}\subseteq\mathbb{R}^{d}$ be a bounded and closed ball of radius $D$ centered at the zero vector. Furthermore, we assume that $\mathcal{K}$ contains the unit ball. Otherwise, we can apply an affine transformation to translate the center point of the convex set to the origin. Let $\mathcal{K^{\prime}}=\{(x,1):x\in\mathcal{K}\}$. For any $\delta\in(0,1)$, let $\mathcal{K}_{\delta}=\{x|\frac{1}{1-\delta}x\in\mathcal{K}\}$ and $\mathcal{K^{\prime}_{\delta}}=\{(x,1):x\in\mathcal{K_{\delta}}\}$, respectively.
@@
In this paper, we consider the bandit optimization $(\mathcal{K},\mathcal{F})$, where $\mathcal{F}$ is the set of $\epsilon$-approximately linear functions $f(x)=\theta^{\top}x+\sigma(x)$ with $\|\theta\|\leq G$ and we assume that $|f(x)|\leq 1,\forall x\in\mathcal{K}.$ Bandit optimization for $\epsilon$-approximately linear functions can be defined as the following statement. For every round $t=1,..,T$, the player selects an action $x_{t}\in\mathcal{K}$, without knowing the loss function in advance. The environment, modeled as an oblivious adversary, chooses a sequence of linear loss vectors $\theta_{1},\theta_{2},...,\theta_{t}$ before the interaction begins and independent of the player’s actions. After selecting $x_{t}$, the adversary chooses a perturbation $\sigma_{t}(x_{t})(|\sigma_{t}(x_{t})|\leq\epsilon)$  [^2] Since we assume that $|f(x)|\leq 1,\forall x\in\mathcal{K}$, we only consider the case where $\epsilon<1$. and the player observes only the scalar loss $f_{t}(x_{t})(=\theta^{\top}_{t}x_{t}+\sigma_{t}(x_{t}))$. The goal of the player is to minimize the regret $\sum_{t=1}^{T}f_{t}(x_{t})-\min_{x\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(x)$.
@@ 4 Main Results
1:$T$, parameters $\eta\in\mathbb{R},\delta\in(0,1)$, $\nu$-normal barrier $\mathcal{R}$ on $con(\mathcal{K})$
2:Initialize: $x^{\prime}_{1}=\arg\min_{x^{\prime}\in\mathcal{K^{\prime}_{\delta}}}{\mathcal{R}(x^{\prime})}$
@@
7:  Update $x^{\prime}_{t+1}=\arg\min\limits_{x^{\prime}\in\mathcal{K^{\prime}_{\delta}}}{\eta\sum_{\tau=1}^{t}g_{\tau}^{\top}x^{\prime}+\mathcal{R}(x^{\prime})}$
@@ 4.1 SCRiBLe with lifting
For the decision set $\mathcal{K}$ with a $\nu$-normal barrier on $con(\mathcal{K})$, where $con(\mathcal{K})=\{\mathbf{0}\}\cup\{(x,b):\frac{x}{b}\in\mathcal{K},x\in\mathbb{R}^{d},b>0\}$, we apply Algorithm 1 to $\epsilon$-approximately linear functions. Recall $\mathcal{K^{\prime}_{\delta}}=\{(x,1):x\in\mathcal{K_{\delta}}\}$.
We simplify SCRiBLe with lifting and increasing learning rates [lee2020bias]. The algorithm performs original SCRiBLe [abernethy2012interior] over a lifted decision set, using a $\nu$-normal barrier $\mathcal{R}$ defined over the $con(\mathcal{K})$ (which always exists) as the regularizer to generate the sequence $x^{\prime}_{1},...,x^{\prime}_{t}$. It set $y^{\prime}_{t}=x^{\prime}_{t}+\mathbf{A}_{t}\mu_{t}$, where $\mathbf{A}_{t}=[\nabla^{2}\mathcal{R}(x^{\prime}_{t})]^{-\frac{1}{2}}$ and $\mu_{t}$ is uniformly sampled at random from $\mathbb{S}_{1}^{d+1}\cap(\mathbf{A}_{t}e_{d+1})^{\perp}$. Since $\mu_{t}$ is orthogonal to $\mathbf{A}_{t}e_{d+1}$, the last coordinate of $\mathbf{A}_{t}\mu_{t}$ is zero, ensuring that $y^{\prime}_{t}=(y_{t},1)$ remains within $\mathcal{K^{\prime}}$. The actual point played is still $y_{t}$. After playing $y_{t}$ and observing $f_{t}(y_{t})(=\theta_{t}^{\top}y_{t}+\sigma_{t}(y_{t}))$, it constructs the loss estimator the same way as SCRiBLe algorithm [abernethy2012interior]: $g_{t}=df_{t}(y_{t})\mathbf{A}_{t}^{-1}\mu_{t}$. Furthermore, it adopts the same update method as FTRL algorithm [hazan2016introduction]: $x^{\prime}_{t+1}=\arg\min\limits_{x^{\prime}\in\mathcal{K^{\prime}}}{\eta\sum_{\tau=1}^{t}g_{\tau}^{\top}x^{\prime}+\mathcal{R}(x^{\prime})}$.
Compared with SCRiBLe with lifting and increasing learning rates [lee2020bias], which focuses on bandit linear optimization, we do not use the increasing learning rates part, but retain the lifting. This preserves its advantages; for instance, the $\nu$-normal barrier $\mathcal{R}$ always exists on $con(\mathcal{K})$ and we can leverage the properties of the $\nu$-normal barrier. When the loss takes the form $f_{t}(y_{t})=\theta_{t}^{\top}y_{t}$ (i.e., when $\epsilon=0$), the analysis by Lee et al. [lee2020bias] shows that the first $d$ coordinates of $g_{t}$ are indeed an unbiased estimator of $\theta_{t}$. However, when $\epsilon\neq 0$ since $g_{t}$ will always be influenced by $\sigma_{t}(y_{t})$, $g_{t}$ is no longer an unbiased estimator of $\theta_{t}$.
@@
The algorithm with parameters $\eta=\frac{\sqrt{\nu\ln\frac{1}{\delta}}}{2d\sqrt{T}}$ and
$\delta=\begin{cases}\frac{1}{T^{2}},&\epsilon=0,\\ \sqrt{\epsilon},&\epsilon\neq 0,\end{cases}$
guarantees the following expected regret bound
$\mathbb{E}\Bigg[\sum_{t=1}^{T}f_{t}(y_{t})-\min_{x\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(x)\Bigg]\leq 4d\sqrt{\nu T\ln\frac{1}{\delta}}+2dT(\nu+2\sqrt{\nu})(\frac{1-\delta}{\delta})\epsilon+\delta GDT+2T\epsilon.$ (2)
@@
The algorithm with parameters $\eta=\frac{\sqrt{\nu\ln\frac{1}{\delta}}}{2d\sqrt{T}}$ and
$\delta=\begin{cases}\frac{1}{T^{2}},&\epsilon=0,\\ \sqrt{\epsilon},&\epsilon\neq 0,\end{cases}$
ensures that with probability at least $1-\gamma$
$\begin{array}[]{l}\sum_{t=1}^{T}f_{t}(y_{t})-\min_{x\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(x)\leq 4d\sqrt{\nu T\ln\frac{1}{\delta}}+2dT(\nu+2\sqrt{\nu})(\frac{1-\delta}{\delta})\epsilon+\delta GDT\\ +2T\epsilon+C(2GD\ln\frac{C}{\gamma}+(1+\epsilon)\sqrt{8T\ln\frac{C}{\gamma}})\end{array}$